# AdaRC: Mitigating Graph Structure Shifts during Test-Time

### v1

@@ 5.3 More discussions
Figure 6: Convergence of AdaRC on Syn-Cora with different learning rates $\eta$.
@@
While AdaRC adapts only the hop-aggregation parameters ${\bm{\gamma}}$ to improve representation quality, other strategies exist, such as adapting the MLP parameters ${\bm{\theta}}$ or both ${\bm{\gamma}}$ and ${\bm{\theta}}$ together. As shown in Figure 6, adapting only ${\bm{\theta}}$ fails to significantly reduce the PIC loss or improve accuracy. Adapting both ${\bm{\gamma}}$ and ${\bm{\theta}}$ minimizes the PIC loss but leads to model forgetting, causing an initial accuracy increase followed by a decline. In contrast, adapting only ${\bm{\gamma}}$ results in smooth loss convergence and stable accuracy, demonstrating that AdaRC effectively adapts to structure shifts without forgetting source graph information. We also compare our proposed PIC loss to other surrogate losses in Appendix C.5. Our PIC loss has better performance under four structure shift scenarios.
@@
AdaRC only introduces two hyperparameters including the learning rate $\eta$ and the number of epochs $T$. In Figure 6, we explore different combinations of them. We observe that AdaRC converges smoothly in just a few epochs, and the final loss and accuracy are quite robust to various choices of the learning rate. Additionally, as discussed in Appendix C.6, we examine the effect of the dimension of hop-aggregation parameters $K$ on AdaRC, and find that it consistently provides stable accuracy gains across a wide range of $K$ values.
@@
We quantify the additional computation time introduced by AdaRC during the test-time. Compared to the standard inference time, AdaRC only adds an extra 11.9% in computation time for each epoch of adaptation. In comparison, GTrans and SOGA adds 486% and 247% in computation time. AdaRC enjoys great efficiency resulting from only updating the hop-aggregation parameters and efficient loss design. Please refer to Appendix C.7 for more details.
@@
Besides GPRGNN, AdaRC is compatible with various GNN architectures, e.g., JKNet [43], APPNP [17], and GCNII [6]. In Appendix C.8, we test the performance of AdaRC with these networks on Syn-Cora. AdaRC consistently improves the accuracy.
@@ 6 Conclusion
In this paper, we explore why generic TTA algorithms perform poorly under structure shifts. Theoretical analysis reveals that attribute structure shifts on graphs bear distinct impact patterns on the GNN performance, where the attribute shifts introduce classifier bias while the structure shifts degrade the node representation quality. Guided by this insight, we propose AdaRC, a plug-and-play TTA framework that restores the node representation quality with convergence guarantee. Extensive experiments consistently and significantly demonstrate the effectiveness of AdaRC.
@@ A.1 Illustration of representation degradation and classifier bias
A.2 Proof of Proposition 3.1 and Corollary 3.2
Proposition 3.1.
@@
When $\gamma\to\infty$, this proposition matches with the results in [25]. [^2] Notice that our notation is slightly different: we use the covariance matrix while they use the square root of it in the multivariate Gaussian distribution.
Corollary 3.2.
@@
A.5 Proof of Proposition 3.5
In this part, instead of treating $\gamma$ as fixed hyperparameter (as in Proposition 3.3 and 3.4), we now consider $\gamma$ as a trainable parameter that can be optimizer on both source and target graphs. We first derive the optimal $\gamma$ for a graph in Lemma A.2
@@
A.6 PIC loss decomposition
@@ Appendix B Convergence analysis
B.1 Convergence of AdaRC
In this section, we give a convergence analysis of our AdaRC framework. For the clarity of theoretical derivation, we first introduce the notation used in our proof.
@@ B.1 Convergence of AdaRC
Figure 8: Computation graph of AdaRC
Figure 8 gives a computation graph of AdaRC.
@@
    to represent the update direction of AdaRC.
Clearly, the convergence of AdaRC depends on the property of the baseline TTA algorithm BaseTTA. In the worst scenario, when the BaseTTA is unreliable and makes completely different predictions in each epoch, the convergence of AdaRC could be challenging. However, in the more general case with mild assumptions on the loss function and baseline TTA algorithm, we show that AdaRC can guarantee to converge. We start our proof by introducing assumptions.
@@
Assumption B.5 is a common assumption in the analysis of convergence [4].
@@
Lemma B.7 gives a general convergence guarantee of noisy gradient descent on smooth functions. Next, in Theorem B.8, we give the convergence analysis of AdaRC.
Theorem B.8 (Convergence of AdaRC).
@@ B.2 Example: linear layer followed by softmax
Lemma B.9 verifies assumption B.1 and B.2: $L_{1}\cdot L_{2}=2\|{\bm{W}}\|_{2}$. It also reveals the benefit of using soft-predictions instead of hard-predictions. Hard predictions can be seen as scaling up ${\bm{W}}$. In this case, the Lipschitz constant will be much larger or even unbounded, which impedes the convergence of AdaRC.
@@ C.1 Effect of attribute shifts and structure shifts
We empirically verify that attribute shifts and structure shifts impact the GNN’s accuracy on target graph in different ways. We use t-SNE to visualize the node representations on CSBM dataset under attribute shifts and structure shifts (homophily shifts). As shown in Figure 9, under attribute shift (c), although the node representations are shifted from the source graph, two classes are still mostly discriminative, which is similar to the case without distribution shifts (c). However, under homophily shift (d), the node representations for two classes mix together. These results match with our theoretical analysis in Propositions 3.3 and 3.4.
@@ C.2 Robustness to noisy prediction
In AdaRC, representation quality and prediction accuracy mutually reinforce each other throughout the adaptation process. A natural question arises: if the model’s predictions contain significant noise before adaptation, can AdaRC still be effective? To address this, we conducted an empirical study on the CSBM dataset with severe homophily shift. We visualize the logits distribution for two classes of nodes in Figure 10.
@@
  • However, as adaptation progresses, AdaRC is still able to gradually refine the node representations and improve accuracy.
Figure 10: AdaRC improves accuracy even when the initial predictions are highly noisy
@@ C.3 Different levels of structure shift
In the main text, we evaluated the performance of AdaRC under both homophily and degree shifts. In this section, we extend our evaluation by testing AdaRC across varying degrees of these structure shifts. For each scenario (e.g., homophily: homo $\to$ hetero, hetero $\to$ homo, and degree: high $\to$ low, low $\to$ high), we manipulate either the homophily or degree of the source graph while keeping the target graph fixed, thereby creating different levels of homophily or degree shifts. The larger the discrepancy between the source and target graphs in terms of homophily or degree, the greater the level of structure shift. For instance, a shift from 0.6 $\to$ 0.2 indicates training a model on a source graph with homophily 0.6 and evaluating it on a target graph with homophily 0.2. By comparison, a shift from 0.8 $\to$ 0.2 represents a more substantial homophily shift.
The results of our experiments are summarized in Tables 3 and 4. Across all four settings, as the magnitude of the structure shift increases, the performance of GNNs trained using ERM declines significantly. However, under all settings, AdaRC consistently improves model performance. For example, in the homo $\to$ hetero setting, when the homophily gap increases from 0.4 (0.6 - 0.2) to 0.6 (0.8 - 0.2), the accuracy of ERM-trained models decreases by over 16%, while the accuracy of models trained with AdaRC declines by less than 2%. This demonstrates that AdaRC effectively mitigates the negative impact of structure shifts on GNNs.

### v2

@@ 5.3 More discussions
Figure 6: Convergence of Matcha on Syn-Cora with different learning rates $\eta$.
@@
While Matcha adapts only the hop-aggregation parameters ${\bm{\gamma}}$ to improve representation quality, other strategies exist, such as adapting the MLP parameters ${\bm{\theta}}$ or both ${\bm{\gamma}}$ and ${\bm{\theta}}$ together. As shown in Figure 6, adapting only ${\bm{\theta}}$ fails to significantly reduce the PIC loss or improve accuracy. Adapting both ${\bm{\gamma}}$ and ${\bm{\theta}}$ minimizes the PIC loss but leads to model forgetting, causing an initial accuracy increase followed by a decline. In contrast, adapting only ${\bm{\gamma}}$ results in smooth loss convergence and stable accuracy, demonstrating that Matcha effectively adapts to structure shifts without forgetting source graph information. We also compare our proposed PIC loss to other surrogate losses in Appendix C.6. Our PIC loss has better performance under four structure shifts.
@@
Matcha only introduces two hyperparameters including the learning rate $\eta$ and the number of epochs $T$. In Figure 6, we explore different combinations of them. We observe that Matcha converges smoothly in just a few epochs, and the final loss and accuracy are quite robust to various choices of the learning rate. Additionally, as discussed in Appendix C.7, we examine the effect of the dimension of hop-aggregation parameters $K$ on Matcha, and find that it consistently provides stable accuracy gains across a wide range of $K$ values.
@@
We quantify the additional computation time introduced by Matcha during the test-time. Compared to the standard inference time, Matcha only adds an extra 11.9% in computation time for each epoch of adaptation. In comparison, GTrans and SOGA adds 486% and 247% in computation time. Matcha enjoys great efficiency resulting from only updating the hop-aggregation parameters and efficient loss design. Please refer to Appendix C.8 for more details.
@@
Besides GPRGNN, Matcha is compatible with various GNN architectures, e.g., JKNet (Xu et al., 2018), APPNP (Klicpera et al., 2019), and GCNII (Chen et al., 2020). In Appendix C.10, we test the performance of Matcha with these networks on Syn-Cora. Matcha consistently improves the accuracy.
@@ 6 Conclusion
In this paper, we explore why generic TTA algorithms perform poorly under structure shifts. Theoretical analysis reveals that attribute structure shifts on graphs bear distinct impact patterns on the GNN performance, where the attribute shifts introduce classifier bias while the structure shifts degrade the node representation quality. Guided by this insight, we propose Matcha, a plug-and-play TTA framework that restores the node representation quality with convergence guarantee. Extensive experiments consistently and significantly demonstrate the effectiveness of Matcha.
Acknowledgment
This work is supported by National Science Foundation under Award No. IIS-2416070, and the U.S. Department of Homeland Security under Grant Award Number 17STQAC00001-08-00. The views and conclusions are those of the authors and should not be interpreted as representing the official policies of the funding agencies or the government.
@@ A.1 Illustration of representation degradation and classifier bias
A.2 Illustration of adapting $\gamma$
In the end of subsection 3.3, we provide two examples to intuitively illustrate why the hop-aggregation parameter $\gamma$ should be adjusted based on the target graph’s degree $d_{T}$ and homophily $h_{T}$. To further demonstrate the effect of adapting $\gamma$, we visualize these examples:
Figure 8: Visualization of effect of structure shifts and adaptation of $\gamma$.
  • Source Graph: We consider a source graph with $\mu_{+}=1$, $\mu_{-}=-1$, $h_{S}=1$, and $d_{S}=10$. In this case, the optimal featurizer assigns equal weight to the node itself and each of its neighbors, resulting in the optimal $\gamma_{S}=10$. Figure 8(a) and (d) show the distribution of node representations on the source graph, where the two classes are well-separated.
  • Homophily Shift: In Figure 8(b), the degree remains unchanged, but the homophily $h_{T}$ decreases to 0.5, meaning each node’s neighbors are equally likely to belong to either class. The neighbors no longer provide reliable information for classification, introducing noise and reducing accuracy to 62%. However, in Figure 8 (c), after adjusting $\gamma_{T}$ to the optimal $d_{T}(2h_{T}-1)=0$, the accuracy improves significantly to 84%.
  • Degree Shift: In Figure 8(e), the homophily remains unchanged, but the degree $d_{T}$ decreases to 1. The original $\gamma$ overemphasizes the neighbors’ representations, placing excessive weight on them, leading to an accuracy drop to 87%. By adjusting $\gamma_{T}$ to the optimal $d_{T}(2h_{T}-1)=1$ in Figure 8(f), the accuracy improves to 92%.
A.3 Proof of Proposition 3.2 and Corollary 3.2
[…]
@@ A.2 Illustration of adapting $\gamma$
When $\gamma\to\infty$, this proposition matches with the results in (Ma et al., 2022). [^2] Notice that our notation is slightly different: we use the covariance matrix while they use the square root of it in the multivariate Gaussian distribution.
[…]
@@
A.6 Proof of Proposition 3.3
In this part, instead of treating $\gamma$ as fixed hyperparameter (as in Proposition 3.2 and 3.2), we now consider $\gamma$ as a trainable parameter that can be optimizer on both source and target graphs. We first derive the optimal $\gamma$ for a graph in Lemma A.2
@@
A.7 Generalization to multi-hop GCNs
So far, our theoretical analysis has focused on single-layer GCNs and CSBM graphs, where homophily and degree are used to parameterize structure shifts in one-hop neighborhoods. In this section, we extend our analysis to more general multi-hop scenarios.
For a graph ${\mathcal{G}}$ and a node $v_{i}$ on the graph, we define its $k$-hop neighbors as follows:
${\mathbb{N}}^{(k)}(v_{i})=\begin{cases}\{v_{i}\},&\text{if }k=0,\\ \{v_{j}:\text{shortest\_distance}(v_{j},v_{i})=k\},&\text{if }k\geq 1.\end{cases}$ (13)
We further define
$d^{(k)}$ $=\frac{1}{N}d_{i}^{(k)},$ where $\text{where }d_{i}^{(k)}$ $=\left|{\mathbb{N}}_{i}^{(k)}\right|,$ (14)
$h^{(k)}$ $=\frac{1}{N}h_{i}^{(k)},$ where $\text{where }h_{i}^{(k)}$ =|$=\frac{|\{v_{j}\in{\mathbb{N}}_{i}^{(k)}:y_{j}=y_{i}\}|}{d_{i}^{(k)}}.$ (15)
For $k=1$, ${\mathbb{N}}_{i}^{(1)}$ corresponds to the standard definition of node neighbors, and $d_{i}^{(1)}$ and $h_{i}^{(1)}$ align with node degree and homophily as defined in the main text. For $k\geq 2$, $d_{i}^{(k)}$ and $h_{i}^{(k)}$ capture more complex structure shifts in the egograph of $v_{i}$. For example:
  • If the 2-hop egograph of $v_{i}$ only contains $d+1$ nodes and all nodes are fully-connected (clustering coefficient is 1), then $d_{i}^{(2)}=0$.
  • If the 2-hop egograph of $v_{i}$ is a tree (clustering coefficient is 0) and each 1-hop neighbor of $v_{i}$ has degree $d$, then $d_{i}^{(2)}=d_{i}^{(1)}\cdot(d-1)$.
This formulation allows us to model higher-order graph metrics, such as cluster coefficients, enabling the analysis of more intricate structure shifts.
Inspired by methods like MixHop (Abu-El-Haija et al., 2019) and GPRGNN (Chien et al., 2021), we consider a $K$-hop GCN with a featurizer defined as:
${\bm{z}}_{i}=\sum_{k=1}^{K}\gamma_{k}\cdot\frac{1}{d_{i}^{(k)}}\sum_{v_{j}\in{\mathbb{N}}^{(k)}(v_{i})}{\bm{x}}_{j},$ (16)
and the linear classifier is still ${\bm{z}}_{i}^{\top}{\bm{w}}+b$.
We denote ${\bm{\gamma}}=[\gamma_{0},\cdots,\gamma_{K}]^{\top}$, and similarly,
${\bm{d}}$ $=[d^{(0)},\cdots,d^{(K)}]^{\top},$ ${\bm{d}}_{i}$ $=[d_{i}^{(0)},\cdots,d_{i}^{(K)}]^{\top},$ (17)
${\bm{h}}$ $=[h^{(0)},\cdots,h^{(K)}]^{\top},$ ${\bm{h}}_{i}$ $=[h_{i}^{(0)},\cdots,h_{i}^{(K)}]^{\top}.$ (18)
Proposition A.3.
For graphs with two classes ${\mathbb{C}}_{+}$ and ${\mathbb{C}}_{-}$, and node attributes ${\bm{x}}_{i}\sim{\mathcal{N}}({\bm{\mu}}_{i},{\bm{I}})$ for each node $v_{i}$, where ${\bm{\mu}}_{i}={\bm{\mu}}_{+}$ for $v_{i}\in{\mathbb{C}}_{+}$ and ${\bm{\mu}}_{i}={\bm{\mu}}_{-}$ for $v_{i}\in{\mathbb{C}}_{-}$, the node representation ${\bm{z}}_{i}$ of node $v_{i}\in{\mathbb{C}}_{+}$ generated by a $K$-hop GCN follows a Gaussian distribution of
${\bm{z}}_{i}\sim{\mathcal{N}}\left(\left({\bm{\gamma}}^{\top}{\bm{h}}_{i}\right)\cdot{\bm{\mu}}_{+}+\left({\bm{\gamma}}^{\top}({\bm{1}}-{\bm{h}}_{i})\right)\cdot{\bm{\mu}}_{-},\left({\bm{\gamma}}^{\top}\operatorname{diag}({\bm{d}}_{i})^{-1}{\bm{\gamma}}\right)\cdot{\bm{I}}\right).$ (19)
When ${\bm{\mu}}_{+}={\bm{\mu}}$, ${\bm{\mu}}_{-}=-{\bm{\mu}}$, and all nodes have the same ${\bm{d}}_{i}$ and ${\bm{h}}_{i}$, the classifier maximizes the expected accuracy when ${\bm{w}}=\operatorname{sign}({\bm{\gamma}}^{\top}(2{\bm{h}}-{\bm{1}}))\cdot\frac{{\bm{\mu}}}{\|{\bm{\mu}}\|_{2}}$ and $b=0$. It gives a linear decision boundary of $\{{\bm{z}}:{\bm{z}}^{\top}{\bm{w}}=0\}$ and the expected accuracy
$\text{Acc}=\Phi\left(\sqrt{\frac{({\bm{\gamma}}^{\top}(2{\bm{h}}-{\bm{1}}))^{2}}{{\bm{\gamma}}^{\top}\operatorname{diag}({\bm{d}}_{i})^{-1}{\bm{\gamma}}}}\cdot\|{\bm{\mu}}\|_{2}\right),$ (20)
where $\Phi$ is the CDF of the standard normal distribution.
Proposition A.3 is a natural extension of Proposition 3.2 and corollary 3.2 from one-layer GCN to multi-hop GCN. Based on this proposition, we can similarly derive conclusions analogous to Propositions 3.3 and 3.4 in the main text.
Finally, we show that the optimal ${\bm{\gamma}}$ should still be adapted based on ${\bm{d}}$ and ${\bm{h}}$, reinforcing the theoretical insights provided earlier.
Lemma A.4.
Under the same setting as Proposition A.3, the optimal ${\bm{\gamma}}$ that maximized the expected accuracy is given by
${\bm{\gamma}}\propto\operatorname{diag}({\bm{d}})(2{\bm{h}}-{\bm{1}})={\bm{d}}\odot(2{\bm{h}}-{\bm{1}}),$ (21)
where $\odot$ is the element-wise multiplication. When $\gamma_{0}=1$, this yields:
$\gamma_{k}=d^{(k)}(2h^{(k)}-1),\quad\forall k=1,\cdots,K,$ (22)
which directly recovers the conclusion of Proposition 3.3.
Proof.
We aim to find ${\bm{\gamma}}$ that maximize the accuracy, equivalently, we optimize the following objective,
${\bm{\gamma}}^{*}=\operatorname*{arg\,max}_{{\bm{\gamma}}}\frac{{\bm{\gamma}}^{\top}(2{\bm{h}}-{\bm{1}})(2{\bm{h}}-{\bm{1}})^{\top}{\bm{\gamma}}}{{\bm{\gamma}}^{\top}\operatorname{diag}({\bm{d}})^{-1}{\bm{\gamma}}}.$ (23)
Let ${\bm{\theta}}=\operatorname{diag}({\bm{d}})^{-\frac{1}{2}}{\bm{\gamma}}$, we solve
${\bm{\theta}}^{*}=\operatorname*{arg\,max}_{{\bm{\theta}}}\frac{{\bm{\theta}}^{\top}\operatorname{diag}({\bm{d}})^{\frac{1}{2}}(2{\bm{h}}-{\bm{1}})(2{\bm{h}}-{\bm{1}})^{\top}\operatorname{diag}({\bm{d}})^{\frac{1}{2}}{\bm{\theta}}}{{\bm{\theta}}^{\top}{\bm{\theta}}}.$ (24)
Therefore, ${\bm{\theta}}^{*}$ is the first eigenvector of $\operatorname{diag}({\bm{d}})^{\frac{1}{2}}(2{\bm{h}}-{\bm{1}})(2{\bm{h}}-{\bm{1}})^{\top}\operatorname{diag}({\bm{d}})^{\frac{1}{2}}$, which is ${\bm{\theta}}^{*}\propto\operatorname{diag}({\bm{d}})^{\frac{1}{2}}(2{\bm{h}}-{\bm{1}})$, and therefore ${\bm{\gamma}}^{*}\propto\operatorname{diag}({\bm{d}})(2{\bm{h}}-{\bm{1}})={\bm{d}}\odot(2{\bm{h}}-{\bm{1}})$.
When $\gamma_{0}^{*}=1$, since $d^{(0)}=1$, $h^{(0)}=1$, we have ${\bm{\gamma}}^{*}={\bm{d}}\odot(2{\bm{h}}-{\bm{1}})$. ∎
A.8 PIC loss decomposition
@@ Appendix B Convergence analysis
B.1 Convergence of Matcha
In this section, we give a convergence analysis of our Matcha framework. For the clarity of theoretical derivation, we first introduce the notation used in our proof.
@@ B.1 Convergence of Matcha
Figure 9: Computation graph of Matcha
Figure 9 gives a computation graph of Matcha.
@@
    to represent the update direction of Matcha.
Clearly, the convergence of Matcha depends on the property of the baseline TTA algorithm BaseTTA. In the worst scenario, when the BaseTTA is unreliable and makes completely different predictions in each epoch, the convergence of Matcha could be challenging. However, in the more general case with mild assumptions on the loss function and baseline TTA algorithm, we show that Matcha can guarantee to converge. We start our proof by introducing assumptions.
@@
Assumption B.5 is a common assumption in the analysis of convergence (Bubeck, 2015).
@@
Lemma B.7 gives a general convergence guarantee of noisy gradient descent on smooth functions. Next, in Theorem B.8, we give the convergence analysis of Matcha.
Theorem B.8 (Convergence of Matcha).
@@ B.2 Example: linear layer followed by softmax
Lemma B.9 verifies assumption B.1 and B.2: $L_{1}\cdot L_{2}=2\|{\bm{W}}\|_{2}$. It also reveals the benefit of using soft-predictions instead of hard-predictions. Hard predictions can be seen as scaling up ${\bm{W}}$. In this case, the Lipschitz constant will be much larger or even unbounded, which impedes the convergence of Matcha.
@@ C.1 Effect of attribute shifts and structure shifts
We empirically verify that attribute shifts and structure shifts impact the GNN’s accuracy on target graph in different ways. We use t-SNE to visualize the node representations on CSBM dataset under attribute shifts and structure shifts (homophily shifts). As shown in Figure 10, under attribute shift (c), although the node representations are shifted from the source graph, two classes are still mostly discriminative, which is similar to the case without distribution shifts (c). However, under homophily shift (d), the node representations for two classes mix together. These results match with our theoretical analysis in Propositions 3.2 and 3.2.
@@ C.2 Robustness to noisy prediction
In Matcha, representation quality and prediction accuracy mutually reinforce each other throughout the adaptation process. A natural question arises: if the model’s predictions contain significant noise before adaptation, can Matcha still be effective? To address this, we conducted an empirical study on the CSBM dataset with severe homophily shift. We visualize the logits distribution for two classes of nodes in Figure 11.
@@
  • However, as adaptation progresses, Matcha is still able to gradually refine the node representations and improve accuracy.
Figure 11: Matcha improves accuracy even when the initial predictions are highly noisy
@@ C.3 Different levels of structure shift
In the main text, we evaluated the performance of Matcha under both homophily and degree shifts. In this section, we extend our evaluation by testing Matcha across varying degrees of these structure shifts. For each scenario (e.g., homophily: homo $\to$ hetero, hetero $\to$ homo, and degree: high $\to$ low, low $\to$ high), we manipulate either the homophily or degree of the source graph while keeping the target graph fixed, thereby creating different levels of homophily or degree shifts. The larger the discrepancy between the source and target graphs in terms of homophily or degree, the greater the level of structure shift. For instance, a shift from 0.6 $\to$ 0.2 indicates training a model on a source graph with homophily 0.6 and evaluating it on a target graph with homophily 0.2. By comparison, a shift from 0.8 $\to$ 0.2 represents a more substantial homophily shift.
The results of our experiments are summarized in Tables 3 and 4. Across all four settings, as the magnitude of the structure shift increases, the performance of GNNs trained using ERM declines significantly. However, under all settings, Matcha consistently improves model performance. For example, in the homo $\to$ hetero setting, when the homophily gap increases from 0.4 (0.6 - 0.2) to 0.6 (0.8 - 0.2), the accuracy of ERM-trained models decreases by over 16%, while the accuracy of models trained with Matcha declines by less than 2%. This demonstrates that Matcha effectively mitigates the negative impact of structure shifts on GNNs.